OpenStudy (anonymous):

Evaluate each expression if x = 14, y = 5, and z = 6. (18-x)+6yz ---- 2 a.94 b.-86 c.364 d.12.5

OpenStudy (anonymous):

Is this your equation? \[18-x + \frac{ 6yz }{ 2 }\]

OpenStudy (anonymous):

yes 6yz is over 2 its a fraction

OpenStudy (anonymous):

\[18-x+\frac{ 6yz }{ 2 }\] \[18-14+\frac{ 6(5)(6) }{ 2 }\] \[4+\frac{ 30(6) }{ 2 }\] \[4+\frac{ 180 }{ 2 }\] Can you take it from here? Just divide 180 / 2 then add 4 to your answer.

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

94??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

\[4+\frac{ 180 }{ 2 }\] \[4+90\] \[94\]

OpenStudy (anonymous):

No problem.